A sum of $1200 is invested at an interest rate of 4.5 % compounded annually. How many years will it take until the sum exceeds $2100?
Answer by jim_thompson5910(35256)

Plug in the given values and use logs to isolate t.

A = P*(1+r/n)^(n*t)
2100 = 1200*(1+0.045/1)^(1*t)
2100 = 1200*(1+0.045)^t
2100 = 1200*(1.045)^t
2100/1200 = 1.045^t
1.75 = 1.045^t
1.045^t = 1.75
log(1.045^t) = log(1.75)
t*log(1.045) = log(1.75)
t = log(1.75)/log(1.045)
t = 12.71366 (this is approximate)

It will take approximately 12.71366 years for the account to have $2100. Round up to the nearest year, 13 years, and you're guaranteed to have more than $2100 in the account.